Dylan has $16000 invested in two accounts; one account pays 7.5% and the other pays 6.5% interest. If he received $1126 in interest in one year, how much does he have interested in each account?
Answer by richwmiller(17219)

Total amount of money invested: $16000
x+y=16000,
Total yearly interest for the two accounts is: $1126
0.075*x+0.065*y=1126
x=16000-y
Substitute for x
0.075*(16000-y)+0.065*y=1126
Multiply out
1200.0-0.075*y+0.065*y=1126
Combine like terms.
-0.01*y=-74.0
Isolate y
y=$7400.00 at 6.5% earns $481.00 interest
x=16000-y
Calculate x
x=$8600.00 at 7.5% earns $645.00 interest
Check
0.075*8600+0.065*7400=1126
645.0+481.0=1126
1126.0=1126
If 1126.0=1126 is TRUE and neither x nor y is negative then all is well
codeint1